OpenStudy (anonymous):

Solve 2e^(-t) - 2te^(-t) = 0

OpenStudy (freckles):

you can factor the e^(-t) out also e^(-t) is never zero you only have to solve 2-2t=0

OpenStudy (anonymous):

So t=1?

OpenStudy (freckles):

\[2 e^{-t}-2t e^{-t} =0 \\ e^{-t}(2-2t)=0 \\ e^{-t} \neq 0 \\ 2-2t=0 \\ 2t=2 \\ t=1 \text{ yep ! }\]

OpenStudy (anonymous):

Yay. Thanks! :D